# Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference

### v1

@@ Appendix A Pseudocode of MJSD
Algorithm 1 One Iteration of MJSD Algorithm
@@
Dataset Task # Test Input Avg. Input Len
ChatGPT-Prompt Instruction 360 25.2
ChatAlpaca Chat 43.081 277.7
CNNDM Summarization 11,490 3,967.1
Appendix B Energy Consumption Measurement
@@
We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption. We use average energy consumption per token to measure energy efficiency. There is a recent study pointing out the measurement error using nvidia-smi [37]. We follow the three principles proposed in [37] to minimize the error.
Appendix C Configuration
@@
The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving [29]. We set the maximum running time to be an hour for each baseline. We use average tokens/second to measure the inference speed and use perplexity (exponentiated average negative log-likelihood) based on the probability of the large model to measure the output quality. Because different methods might finish different numbers of inputs, we only calculate the perplexity of the first $M$ inputs, where $M$ is the number of inputs finished by greedy decoding. We use average energy consumption per token to measure energy efficiency. The details of energy measurement are illustrated in the Appendix.
Appendix D Hyper-parameter Details
@@
In the experiments, we follow the default settings in [2] to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
@@
Table 5: Hyper-parameters of different methods for different models and datasets. CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
Llama,CP OPT,CP Llama,CA OPT,CA Llama,D OPT,CD
speculative step len $\gamma$ 4 4 4 4 4 4
Spectr step len $\gamma$ 4 4 4 4 4 4
num seq $m$ 4 4 2 2 4 2
BiLD step len $\gamma$ 10 10 10 10 10 10
fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3
rollback thres $\tau_{2}$ 2 2 1 2 3 2
MJSD step len $\gamma$ 4 4 4 4 4 4
num beams 8 8 8 8 8 8
acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1
Appendix E License of Datasets and Models
@@
Appendix F More Discussion on Multi-Token Joint Greedy Decoding
It is not trivial to prove that MJGD is better than vanilla greedy decoding. We might tend to show that given the same $x_{1:t}$ and $prefix$, the $\gamma_{i}$ tokens generated by MJGD has higher likelihood than the $\gamma_{i}$ tokens generated by calling greedy decoding $\gamma_{i}$ times. However, at step $i$ ($\forall i\geq 2$), the previously generated tokens $x_{1:t}$ are different for the two decoding algorithms, making it difficult to compare the overall perplexity.
To explain why MJGD is more effective, let us consider an optimal algorithm that generates next $\gamma_{i}$ tokens at one step based on a score function:
$s^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+1:N}|x_{1:t})$ (3)
which is the log-likelihood of the optimal future sequence with prefix $x_{1:t}$ and the generated $\gamma_{i}$ tokens. The optimal greedy decoding chooses $x_{t+1:t+\gamma_{i}}$ that maximizes the score function and is guaranteed to return the sequence with the optimal perplexity.
Both MJGD and vanilla greedy decoding can be viewed as approximations to the optimal greedy decoding by replacing $s^{*}(\cdot)$ with $s_{MJGD{}}$ or $s_{GD}$ that estimates the values of $s^{*}(\cdot)$. Specifically, since MJGD selects the $\gamma_{i}$ tokens with the largest joint likelihood, we have $s_{MJGD}(\cdot)=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{1}$ where $c_{1}$ is an arbitrary constant. Notice that adding a constant to $s(\cdot)$ changes the error of the estimation but does not change the behavior of the algorithm.
On the other hand, greedy decoding generates the $\gamma_{i}$ tokens one by one. At the last token, greedy decoding selects the token with the largest $p(x_{t+\gamma_{i}}|x_{1:t+\gamma_{i}-1})$, which is equivalent to selecting the tokens with the largest $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ provided that $x_{t+1:t+\gamma_{i}-1}$ are the same tokens generated by previous steps. So for greedy decoding, $s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$ ($c_{2}$ is also an arbitrary constant) if and only if $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$ for $i=1,\ldots,\gamma_{i}-1$. Otherwise, the score of $x_{t+1:t+\gamma_{i}}$ has to be smaller than any possible value of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$. Notice that the lower bound of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})$ is $\infty$- ∞, so $s_{GD}(x_{t+1:t+\gamma_{i}})=-\infty$ if and only if $\exists j=1,\ldots,\gamma_{i}-1$ such that $x_{t+j}\neq\arg\max_{x}p(x|x_{1:t+i-1})$.
Let $c^{*}=\arg\min_{c_{2}}\sum_{x_{t+1:t+\gamma_{i}}}|s_{GD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|$. By setting $c_{1}=c^{*}$, for $x_{t+1},\ldots,x_{t+\gamma_{i}-1}$ satisfying that $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$, $\forall i=1,\ldots,\gamma_{i}-1$, we have
$s_{MJGD}(x_{t+1:t+\gamma_{i}})=s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}|x_{1:t}})+c^{*}$
Otherwise, we have
$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|=|c^{*}-h^{*}(x_{t+1:t+\gamma_{i}})|\leq|-\inf-s^{*}(x_{t+1:t+\gamma_{i}})|$
where $h^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+\gamma_{i}+1:N}|x_{1:t+\gamma_{i}},prefix)$. Therefore,
$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|\leq|s_{GD}(x_{1:K})-s^{*}(x_{t+1:t+\gamma_{i}})|$ (4)
The derivation above provides a heuristic explanation to show that MJGD is a closer approximation to the optimal greedy decoding than vanilla greedy decoding. Hence, its output perplexity is generally better.

### v2

@@ Appendix B Pseudocode of MJSD
Algorithm 1 One Iteration of MTAD Algorithm
@@
Dataset Task Avg. Input Len
ChatGPT-Prompt Instruction 25.2
ChatAlpaca Chat 277.7
CNNDM Summarization 3,967.1
Spider Text-to-SQL 347.68
MT-Bench Various [^1] The tasks of MT-Bench cover humanities, extraction, roleplay, math, coding, reasoning, stem, writing, and STEM. N/A [^2] MT-Bench contains multi-turn tasks where the input includes the responses of LLMs, so the input length is not fixed.
Appendix C Additional Experiments
C.1 Resutlts with OPT-30B and Llama-2-70B
Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B). Table 9 shows the average performances across all datasets. MTAD always achieves the lowest perplexity and the best efficiency.
Table 9: Inference efficiency and output perplexity of different methods with OPT (350M,30B) and Llama-2-Chat (7B,70B). The mean and standard deviation are computed across all datasets. Bold numbers mark the best result, underlined numbers mark the second best.
[…]
energy (J/token) $\downarrow$ 138$\pm$87.7 142$\pm$99.7 122$\pm$66.4 125$\pm$65.4 119$\pm$67.7
perplexity $\downarrow$ 1.77$\pm$0.22 1.69$\pm$0.25 1.73$\pm$0.24 1.73$\pm$0.24 1.52$\pm$0.19
OPT speed (token/s) $\uparrow$ 15.3$\pm$1.64 14.5$\pm$1.96 17.0$\pm$4.14 17.4$\pm$4.00 19.5$\pm$4.11
energy (J/token) $\downarrow$ 72.4$\pm$11.5 79.6$\pm$3.03 68.2$\pm$16.7 62.4$\pm$10.3 60.0$\pm$12.8
perplexity $\downarrow$ 4.74$\pm$1.96 3.50$\pm$1.42 4.55$\pm$1.93 4.49$\pm$ 1.95 2.74$\pm$0.87
[…]
C.2 Ablation study of top-k and top-p sampling
Here we conduct an ablation study to show how the value of $k$ and $p$ in top-k and top-p warping affects our method. Table 10 shows the results. We can see that when changing the value of $k$ and $p$, MTAD consistently achieves significantly better performances.
Table 10: Ablation study of $k$ and $p$ in top-k and top-p sampling
K P Greedy Speculative MJSD
PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
C.3 Additional Experiments on CNNDM and Spider
We also report the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset. Table 11 and Table 12 show the results. We can see that MTAD consistenly achieve better effectiveness as well as faster decoding speed.
Table 11: Comparison of ROUGE-L Scores and Tokens per Second under Different Fine-Tuning Conditions on CNNDM
No Fine-Tune Fine-Tune 68M Fine-Tune Both
Method ROUGE-L Tokens/sec ROUGE-L Tokens/sec ROUGE-L Tokens/sec
Speculative 0.114 37.7 0.114 20.4 0.164 24.3
MJSD 0.118 44.2 0.121 25.0 0.168 27.1
Table 12: Comparison of Execution Accuracy (EA) and Tokens per Second under Different Fine-Tuning Conditions on Spider
No Fine-Tune Fine-Tune 68M Fine-Tune Both
Method EA Tokens/sec EA Tokens/sec EA Tokens/sec
Speculative 11.5 28.5 11.5 27.8 16.3 25.6
MJSD 13.0 30.3 14.8 32.3 18.3 29.4
Appendix D Energy Consumption Measurement
@@
We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption. We use average energy consumption per token to measure energy efficiency. There is a recent study pointing out the measurement error using nvidia-smi (Yang et al., 2023). We follow the three principles proposed in (Yang et al., 2023) to minimize the error.
Appendix E Configuration
@@
The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving (Schuster et al., 2022). We set the maximum running time to be an hour for each baseline. We use average tokens/second to measure the inference speed and use perplexity (exponentiated average negative log-likelihood) based on the probability of the large model to measure the output quality. Because different methods might finish different numbers of inputs, we only calculate the perplexity of the first $M$ inputs, where $M$ is the number of inputs finished by greedy decoding. We use average energy consumption per token to measure energy efficiency. The details of energy measurement are illustrated in the Appendix.
Appendix F Hyper-parameter Details
@@
In the experiments, we follow the default settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
@@
Table 13: Hyper-parameters of different methods for different models and datasets. L: Llama, O:OPT, CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
L,CP O,CP L,CA O,CA L,CD O,CD L,SP O,SP L,MT O,MT
speculative step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
Spectr&SpecInfer step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
num seq $m$ 4 4 2 2 4 2 2 2 2 2
BiLD step len $\gamma$ 10 10 10 10 10 10 10 10 10 10
fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3 0.9 0.9 0.9 0.9
rollback thres $\tau_{2}$ 2 2 1 2 3 2 1 1 1 1
MTAD step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
num beams 8 8 8 8 8 8 8 8 8 8
acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1 0.9 0.1 0.9 0.1
Appendix G License of Datasets and Models